There are 2 bonds in the portfolio. Their current prices, interest rates, and durations are listed below. What is the change in the portfolio value due to a 1% drop in the interest rate? Bond A B Market price 800 1,200 Interest rate 7% 5% Duration (year) 6 4
Answer #1

Current portfolio value = MV A + MV B = 800+1200 = 2000

change in MV A = -modified duration*bond price*yield change

= -6*800*(-.01) =48

change in MV B = -modified duration*bond price*yield change

= -4*1200*(-.01) =48

Change in MV of portfolio = change in MV A+change in MV B = 48+48 = 96